OpenStudy (anonymous):

Find the difference negative two-sixths minus negative seven-ninths. Be sure to reduce to simplest form. positive four-ninths negative four-ninths positive five-thirds negative five-thirds

OpenStudy (anonymous):

neg two/sixth - neg seven/ninths = pos 0.444444444 ( pos four/ninths)

OpenStudy (anonymous):

-2/6-(-7/9) =-1/3+7/9 =-3/9+7/9 =4/9